Obesity is defined as a body mass index (BMI)of30kg/m 2 or more. A 95% confidence interval for the percentage of U.S. adults aged 20 years and over who were obese was found to be 22.4 to 23.5%. What was the sample size?

The answer should be 318, but I don't know how to work it out.
Answer #1

The formula for CI is,

media%2Fa28%2Fa284dadf-d772-4329-a1f7-db

So, lower limit = 0.224 = p-z*\sqrt{\frac{p(1-p)}{n}}

and upper limit = 0.235 = p+z*\sqrt{\frac{p(1-p)}{n}}

Now, z score for 95% confidence level is, 1.96

and lower limit + upper limit = 2p = 0.224+0.235 = 0.459

=> p = 0.2295

Putting p and z in lower limit we get,

0.224 = 0.2295-1.96*\sqrt{\frac{0.2295(1-0.2295)}{n}}

=> (0.224-0.2295)/1.96 = -\sqrt{\frac{0.2295(1-0.2295)}{n}}

=> 0.0028 = \sqrt{\frac{0.1768}{n}}

=> n= 0.1768/0.00282=